Chamberlain Enterprises Inc. reported the following receivables in its December 31, 2018, year-end balance sheet: Current assets: Accounts receivable, net of $24,000 in allowance for uncollectible accounts $ 218,000 Interest receivable 6,800 Notes receivable 260,000 Additional Information: The notes receivable account consists of two notes, a $60,000 note and a $200,000 note. The $60,000 note is dated October 31, 2018, with principal and interest payable on October 31, 2019. The $200,000 note is dated June 30, 2018, with principal and 6% interest payable on June 30, 2019. During 2019, sales revenue totaled $1,340,000, $1,280,000 cash was collected from customers, and $22,000 in accounts receivable were written off. All sales are made on a credit basis. Bad debt expense is recorded at year-end by adjusting the allowance account to an amount equal to 10% of year-end accounts receivable. On March 31, 2019, the $200,000 note receivable was discounted at the Bank of Commerce. The bank's discount rate is 8%. Chamberlain accounts for the discounting as a sale. Required: 1. Not including sales revenue, what revenue and expense amounts related to receivables will appear in Chamberlain’s 2019 income statement? 2. & 3. What amounts will appear in the 2019 year-end balance sheet for accounts receivable and Calculate the receivables turnover ratio for 2019.

  • Req 1
  • Req 2 and 3

Not including sales revenue, what revenue and expense amounts related to receivables will appear in Chamberlain’s 2019 income statement?

Revenue:
Interest revenue
Expense:
Bad debt expense
Answer #1
Requirement 1:The following Revenue and expenses will occur in 2019 income statement
1 Revenue
$200,000*6%*3/12 =$3,000
$60,000*8%*10/12 =$4,000
Total revenue =$7,000
2 Accounts receivable
Beginning accounts receivable(Gross) =$242,000($218,000+$24,000)
Add:Credit sales =$1,340,000
Less:Write-offs =$(22,000)
Less:Cash collection =(1,280,000)
Ending Acconts Receivable =$280,000
Allowance for uncollectible accounts
Beginning allowance =$24,000
Add:Bad debt expenses(Balancing figure) =$26,000(24,000-22,000-28,000)
Less:Write-offs =$(22,000)
Ending allowance =$28,000(280,000*10%)
3 Loss on sale of Notes receivable =$200,000*6%*9/12 =$9,000
4 Opening Notes receivable =$200,000
Add:Interest on maturity =$12,000($200,000*6%)
Maturity value =$212,000
Less:Discount = $(4,240) ($212,000*8%*3/12)
Cash Proceeds =$207,760
5 Carrying value of Notes Receivable= $209,000($200,000+9,000)
Less:Cash Proceeds =$(207,760)
Loss on sale of notes receivable =$1,240
Req 2 Accounts Receivable =$280,000
Less:Allowance for uncollectible accounts =$28,000
Net accounts receivable for 2019 =$252,000
Req 3 Accounts Receivable Ratio =Net credit sales / Average Accounts receivable
Average Accounts receivable =($218,000+$252,000) / 2 =$235,000
Accounts Receivable Ratio =$1,340,000 / $235,000 =5.7 times